Exercise 1-15 (Algo) Traditional and Contribution Format Income Statements [LO1-6]

The Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter ended March 31:

Amount
Sales $ 946,000
Selling price per pair of skis $ 430
Variable selling expense per pair of skis $ 49
Variable administrative expense per pair of skis $ 18
Total fixed selling expense $ 130,000
Total fixed administrative expense $ 105,000
Beginning merchandise inventory $ 75,000
Ending merchandise inventory $ 120,000
Merchandise purchases $ 315,000

Required:

1. Prepare a traditional income statement for the quarter ended March 31.

2. Prepare a contribution format income statement for the quarter ended March 31.

3. What was the contribution margin per unit?
Answer #1

1) Cost of goods sold = 75000+315000-120000 = 270000

The Alpine House, Inc.
Mar 31
Sales revenue 946000
Cost of goods sold 270000
Gross profit 676000
Selling and administrative expense
Selling expense (2200*49+130000) 237800
Administrative expense (2200*18+105000) 144600
Total Selling and administrative expense 382400
Net income 293600

2) Contribution margin income statement

The Alpine House, Inc.
Mar 31
Sales revenue 946000
Variable cost
Cost of goods sold 270000
Selling expense 107800
Administrative expense 39600
Total Variable cost 417400
Contribution margin 528600
Fixed cost
Selling expense 130000
Administrative expense 105000
Total Fixed cost 235000
Net income 293600

3) Contribution margin per unit = 430-(270000/2200+49+18) = 240.27